find one of the legs for the right-angle triangle in which it hypotenuse is 10 5m and other leg is 7m


1
Expert's answer
2021-09-29T00:12:01-0400

H2= h2+b2

b2=H2-h2

=10.52-72

=110.25-49

b2=61.25

b= (61.25)1/2

b=7.8m